sam needs to get carpet for two rooms of his house. Estimate the number of square feet in the two rooms if one room measure 10 and 1/2 feet by 9 and 3/4 feet; and the other room measures 19 and 1/4 feet by 18 and 1/2 feet.
Answer by checkley71(8403)

10.5*9.75=102.375
19.25*18.5=356.125
102.375+356.125=458.5 square feet of carpet will be required.